Measurements of scientific systems are always subject to variation, some more than others. There are many structures for measurement error, and statisticians spend a great deal of time modeling these errors. Suppose the measurement error X of a certain physical quantity is decided by the density function

f(x)= {k(3-x^2), -1<x<1,
.........0,(included on the bracket) elsewhere

(a) Determine k that renders f(x) a valid density function. (Fraction form)
(b) Find the probability that a random error is more than 0. (Fraction form)
(c) Find the probability that a random error is within -1/2 and 1/2. (Fraction form)
1
Expert's answer
2020-12-13T18:40:58-0500

A) "\int^1_{-1}f(x)dx=1"

"\int^1_{-1}k(3-x^2)dx=(3kx-kx^3/3)|_{-1}^1=3k-k/3+3k-k/3=6k-2k/3=18k/3-2k/3=16k/3=1"

k=3/16

B)"\int_0^1( 3(3-x^2)/16) dx=(9x/16-x^3/16)|_0^1=9/16-1/16=8/16=1/2"

C)"\int_{-1/2}^{1/2}(3 (3-x^2)/16) dx=(9x/16-x^3/16)|_{-1/2}^{1/2}=9/32-1/128+9/32-1/128=18/32-2/128=9/16-1/64=36/64-1/64=35/64"